# Sliced-Wasserstein Estimation with Spherical Harmonics as Control Variates

### v1

@@ B.4 Kernel Support Vector Machines and Sliced-Wasserstein Distances
Kernel Support Vector Machines. Kernel Support Vector Machines (SVMs) are a powerful class of supervised learning algorithms used for classification and regression tasks. At their core, SVMs aim to find a hyperplane that best separates the data into different classes while maximizing the margin between these classes. However, in cases where the classes are not linearly separable, Kernel SVMs come into play. The key idea behind Kernel SVMs is to map the input data $x\in\mathbb{R}^{q}$ into a higher-dimensional space via $\phi(x)$, where a linear separation might be achievable. The kernel function $k(\cdot,\cdot)$ computes the similarity between pairs of data points in the higher-dimensional space without explicitly calculating the transformed data. Given training data $x_{i}\in\mathbb{R}^{q}$ with $i=1,\ldots,n$ divided into two classes and a vector of labels $y\in\{-1;+1\}$, the goal is to find the parameters $\omega\in\mathbb{R}^{q}$ and $b\in\mathbb{R}$ such that the prediction given by $\mathrm{sign}(\omega^{\top}\phi(x_{i})+b)$ is correct for most samples.
@@
6:    ${\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}^{(i,j)}$
7:    $K[i,j]=\exp(-\gamma{\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\pgfsys@color@rgb@stroke{0.00390625}{0.44921875}{0.69921875}\pgfsys@color@rgb@fill{0.00390625}{0.44921875}{0.69921875}\text{MC}}_{n})$
@@
6:    ${\color[rgb]{0.80078125,0.46875,0.73828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.46875,0.73828125}\pgfsys@color@rgb@stroke{0.80078125}{0.46875}{0.73828125}\pgfsys@color@rgb@fill{0.80078125}{0.46875}{0.73828125}\text{SHCV}}_{n}=w_{n}^{\top}f_{n}^{(i,j)}$
7:    $K[i,j]=\exp(-\gamma{\color[rgb]{0.80078125,0.46875,0.73828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.46875,0.73828125}\pgfsys@color@rgb@stroke{0.80078125}{0.46875}{0.73828125}\pgfsys@color@rgb@fill{0.80078125}{0.46875}{0.73828125}\text{SHCV}}_{n})$
@@ D.2 Number of Spherical Harmonics
Figure 8(a) presents the evolution of the number $N_{\ell}^{d}$ of spherical harmonics of degree $\ell$ in dimension $d$ with $d\in[3,10]$ and $\ell\in[1,10]$ while Figures 8(b) and 8(c) report the evolution of the cumulative sum $\sum_{\ell=1}^{L}N_{\ell}^{d}$ and even cumulative sum $\sum_{\ell=1}^{L}N_{2\ell}^{d}$ for maximum degree $L\in[1,10]$ and $2L\in\{2;4;6;8;10;12\}$ respectively.
@@ D.4 Simple Case Study for Discrete Measures and Gaussian Measures
Recall that $\operatorname{SHCV}_{n,L}^{p}(\mu_{2},\nu_{2})=\mathrm{I}_{n}^{\mathrm{ols}}$ where $(\mathrm{I}_{n}^{\mathrm{ols}},\beta_{n})\in\arg\min_{\alpha,\beta}\|f_{n}-\alpha\mathds{1}_{n}-\Phi\beta\|_{2}^{2}$. Figure 9(a) reports the evolution of $\mathrm{W}_{2}^{2}(\theta_{\sharp}^{\star}\mu_{2},\theta_{\sharp}^{\star}\nu_{2})$ (black line) and its polynomial approximation $\widehat{f}_{2L}(\theta)$ with different degrees of harmonics $2L\in\{2;4;8;10\}$. Figure 9(b) displays the approximation error, i.e., the difference $\varepsilon_{n}$ between $\mathrm{W}_{2}^{2}(\theta_{\sharp}^{\star}\mu_{2},\theta_{\sharp}^{\star}\nu_{2})$ and its polynomial approximation $\widehat{f}_{2L}(\theta)$.

### v2

@@ B.4 Kernel Support Vector Machines and Sliced-Wasserstein Distances
Kernel Support Vector Machines. Kernel Support Vector Machines (SVMs) are a powerful class of supervised learning algorithms used for classification and regression tasks. At their core, SVMs aim to find a hyperplane that best separates the data into different classes while maximizing the margin between these classes. However, in cases where the classes are not linearly separable, Kernel SVMs come into play. The key idea behind Kernel SVMs is to map the input data $x\in\mathbb{R}^{q}$ into a higher-dimensional space via $\phi(x)$, where a linear separation might be achievable (see Figure 7 below).
Figure 7: Visualization of Kernel transformation.
The kernel function $k(\cdot,\cdot)$ computes the similarity between pairs of data points in the higher-dimensional space without explicitly calculating the transformed data. Given training data $x_{i}\in\mathbb{R}^{q}$ with $i=1,\ldots,n$ divided into two classes and a vector of labels $y\in\{-1;+1\}$, the goal is to find the parameters $\omega\in\mathbb{R}^{q}$ and $b\in\mathbb{R}$ such that the prediction given by $\mathrm{sign}(\omega^{\top}\phi(x_{i})+b)$ is correct for most samples.
@@
Figure 8: Comparison of different linear SVM classifiers on a 2D projection of the iris dataset.
@@
6:    ${\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n}=(\mathds{1}_{n}/n)^{\top}f_{n}^{(i,j)}$
7:    $K[i,j]=\exp(-\gamma{\color[rgb]{0.00390625,0.44921875,0.69921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.44921875,0.69921875}\text{MC}}_{n})$
@@
6:    ${\color[rgb]{0.80078125,0.46875,0.73828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.46875,0.73828125}\text{SHCV}}_{n}=w_{n}^{\top}f_{n}^{(i,j)}$
7:    $K[i,j]=\exp(-\gamma{\color[rgb]{0.80078125,0.46875,0.73828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.46875,0.73828125}\text{SHCV}}_{n})$
@@
As mentioned in Remark 5, computational benefits occur when $K>1$ integrals need to be computed since the SHCV time complexity scales as $\mathcal{O}(Kn\omega_{f}+\omega(\Phi))$ compared to $\mathcal{O}(Kn\omega_{f})$ for the standard Monte Carlo estimate when dealing with $K$ integrals. Thus, when $K$ is large, the extra term $\omega(\Phi)$ becomes negligible. In the experiments we are dealing with images and probability measures $\mu_{1},\ldots,\mu_{N_{X}},\nu_{1},\ldots,\nu_{N_{Y}}$ in dimension $d=2$ with a number of Monte Carlo projections $n\in\{25;50;75;100\}$. The total number of integrals to be computed are:
• $K_{XX}=\binom{N_{X}}{2}=1437\times 1436/2=1,024,581$ for the ’train/train’ kernel matrix.
• $K_{XY}=N_{X}\times N_{Y}=1437\times 360=517,320$ for the ’train/test’ kernel matrix.
Below, we compare the time (in $\mu s$) to compute the evaluations $f_{n}^{(i,j)}=(f_{\mu_{i},\nu_{j}}^{(2)}(\theta_{k}))_{k=1}^{n}$ for a single pair $(i,j)$ vs. the time to compute the SHCV weights $w_{n}$. These weights need to be computed only once for all pairs of images. Interestingly, this extra computing time is of the same order are the time required for the evaluations of the integrands so that in the end it becomes negligible as the SHCV time complexity scales as $\mathcal{O}(Kn\omega_{f}+\omega(\Phi))$ compared to $\mathcal{O}(Kn\omega_{f})$ for the standard Monte Carlo estimate.
MC projections $n$ evaluations $f_{n}^{(i,j)}$ in $\mathcal{O}(n\omega_{f})$ SHCV weights $w_{n}$ in $\mathcal{O}(\omega(\Phi))$
25 $422\pm 36$ $207\pm 243$
50 $700\pm 16$ $479\pm 516$
75 $945\pm 17$ $761\pm 559$
100 $1182\pm 20$ $1057\pm 654$
Table 8: Comparison of time computations (in $\mu s$) for the evaluations of the integrands vs. SHCV weights.
@@ D.2 Number of Spherical Harmonics
Figure 9(a) presents the evolution of the number $N_{\ell}^{d}$ of spherical harmonics of degree $\ell$ in dimension $d$ with $d\in[3,10]$ and $\ell\in[1,10]$ while Figures 9(b) and 9(c) report the evolution of the cumulative sum $\sum_{\ell=1}^{L}N_{\ell}^{d}$ and even cumulative sum $\sum_{\ell=1}^{L}N_{2\ell}^{d}$ for maximum degree $L\in[1,10]$ and $2L\in\{2;4;6;8;10;12\}$ respectively.
@@ D.4 Simple Case Study for Discrete Measures and Gaussian Measures
Recall that $\operatorname{SHCV}_{n,L}^{p}(\mu_{2},\nu_{2})=\mathrm{I}_{n}^{\mathrm{ols}}$ where $(\mathrm{I}_{n}^{\mathrm{ols}},\beta_{n})\in\arg\min_{\alpha,\beta}\|f_{n}-\alpha\mathds{1}_{n}-\Phi\beta\|_{2}^{2}$.
Figure 10(a) reports the evolution of $\mathrm{W}_{2}^{2}(\theta_{\sharp}^{\star}\mu_{2},\theta_{\sharp}^{\star}\nu_{2})$ (black line) and its polynomial approximation $\widehat{f}_{2L}(\theta)$ with different degrees of harmonics $2L\in\{2;4;8;10\}$. Figure 10(b) displays the approximation error, i.e., the difference $\varepsilon_{n}$ between $\mathrm{W}_{2}^{2}(\theta_{\sharp}^{\star}\mu_{2},\theta_{\sharp}^{\star}\nu_{2})$ and its polynomial approximation $\widehat{f}_{2L}(\theta)$.